OpenStudy (anonymous):

Help me please... your given that: u(base 1)=1 and u(base n+1)=u(base n)/1+u(base n) Find the values of U (base 2), U(base 3) and U (base 4). give your answers as a fraction and please be so nice to explain it

OpenStudy (amistre64):

that looks like recursion formulas

OpenStudy (amistre64):

\[U_{n+1} = \frac{U_{n}}{1+U_{n}}\]

OpenStudy (amistre64):

U{1} = 1 so.... U{2} = 1/(1+1) = 1/2 U{3} = (1/2)/(1+(1/2)) = 1/3 U{4} = [1/3]/ [1+(1/3)] = 1/4

OpenStudy (amistre64):

you just keep reiterating the past amount to find the present amount

OpenStudy (amistre64):

'btm' works with themand his material calls them difference equations..... which seems an archaic term

OpenStudy (anonymous):

thank you very much medal earned :)